OpenStudy (anonymous):

Really need help with 5 questions please!

OpenStudy (anonymous):

Are you busy?

OpenStudy (dtan5457):

its okay ask away

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What is the solution to the equation? 8x + 4 = –20

OpenStudy (anonymous):

8x+4=-20

OpenStudy (anonymous):

isolate x by putting x on one side of the equation and everything else on the other side.

OpenStudy (dtan5457):

if you understood the work i showed before this should be done with ease.

OpenStudy (dtan5457):

8x+4=-20 subtract the 4, divide by 8 to get x alone

OpenStudy (dtan5457):

try it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=2?

OpenStudy (dtan5457):

8x=-24 x=? its not 2

OpenStudy (anonymous):

so 8 times what = -24?

OpenStudy (anonymous):

thats the question?

OpenStudy (dtan5457):